# SCI: An Equilibrium for Signal Intelligence

### v1

@@ Abstract
We present SCI, a closed-loop, control-theoretic framework that models interpretability as a regulated state. SCI formalizes the interpretive error $\Delta SP$ and actively drives $SP(t)\in[0,1]$ (“Surgical Precision”) toward a target via a projected update on $\Theta$ under a human-gain budget. The framework operates through three coordinated components: (1) reliability-weighted, multiscale features $P(t,s)$; (2) a knowledge-guided interpreter $\psi_{\Theta}$ that emits traceable markers and rationales; and (3) a Lyapunov-guided controller equipped with rollback, trust-region safeguards, and a descent condition. Across biomedical (EEG/ECG/ICU), industrial (bearings/tool wear), and environmental (climate/seismic) domains, SCI reduces interpretive error by 25–42% (mean 38%, 95% CI 22–43%) relative to static explainers while maintaining AUC/F1 within $\pm$1–2%. SCI also reduces $SP$ variance from 0.030 to 0.011, indicating substantially more stable explanations. Modeling interpretability as a control objective yields steadier, faster-recovering, and more trustworthy interpretive behavior across diverse signal regimes.
@@ 1 Introduction
In safety-critical settings, the usefulness of an alert depends as much on its rationale as on its score. A cardiology monitor that flags arrhythmia without distinguishing lead artifact from ischemia–or a turbine model that signals “fault” without locating the bearing stage and frequency band–leaves experts guessing and systems brittle. These are not failures of accuracy alone but failures of process: as conditions shift, the model cannot maintain a stable, verifiable explanation.
On the acronym “SCI”.
@@
SCI was conceived to close this gap. Earlier iterations laid the groundwork. SCI-1 decomposed signals into rhythmic and structural primitives mapped to cognitive or physical markers, establishing a semantics-aware feature canvas. SCI-2 introduced Surgical Precision (SPSP)–a quantitative clarity metric–and framed the notion of interpretive equilibrium sustained by feedback correction.
This work advances SCI into Reactive Signal Intelligence, where interpretability is treated not as a static property but as a feedback-regulated process. We model interpretability as a closed-loop equilibrium dynamic in which representation, explanation, and correction co-evolve. Within this control-theoretic formulation, interpretation becomes a quantifiable control objective: the system continuously minimizes an interpretive discrepancy $\Delta SP$ to align internal reasoning with domain reality.
SCI instantiates this dynamic view through three integrated components:
@@
  • A Lyapunov-guided adaptive controller with a human-feedback gain budget $\lambda_{h}$, descent conditions, and stability safeguards (with $\lambda_{h}<\mu/(Uc)$ and trust-region/rollback, $V=\tfrac{1}{2}(\Delta SP)^{2}$ decreases monotonically up to bounded noise; see §5.4). This reframes transparency as a control problem that sustains equilibrium-maintaining both performance and clarity under sensor drift, nonstationarity, and bounded perturbations (adversarial or human).
@@
  • Stability and Evidence. We provide theoretical stability conditions and empirical validation, demonstrating $25--40\%$ reductions in $\Delta SP$ across biomedical, industrial, and environmental domains while preserving task performance and reducing explanation variance.
@@
  • Section 2 motivates reactivity with domain examples.
  • Section 3 states objectives and delineates contributions.
  • Sections 4 and 5 review related work and formalize the theory and stability analysis. Section 6 details the architecture and feedback mechanisms.
  • Section 7 evaluates SCI across EEG/ECG/ICU, bearings/tool wear, and climate/seismic signals.
  • Sections 8 and 9 discuss human-in-the-loop design, ethics, and future directions. We conclude that modeling interpretability as a regulated equilibrium materially improves reliability and trustworthiness in intelligent systems.
@@
SCI in 60 seconds. What: Treat explanation quality as a signal $SP(t)$ with target $SP^{*}$. How: When $|\Delta SP|{=}|SP^{*}-SP|$ is large, update $\Theta$ via $\Theta\leftarrow\mathrm{Proj}_{\mathcal{C}}\!\big[\Theta+\eta(\Delta SP\,\nabla_{\Theta}SP+\lambda_{h}u_{h})\big]$, bounded by rollback/trust-region and a human-gain budget $\lambda_{h}<\mu/(Uc)$. Why it works: With these bounds, the Lyapunov energy $V=\tfrac{1}{2}(\Delta SP)^{2}$ decreases (up to noise), stabilizing rationales while preserving task accuracy. Payoff: In 7 streaming tasks, $\Delta SP\downarrow$ 25–42% with steady AUC/F1 and $\sim$63% lower $SP$ variance.
@@ 2.1 Real-world failure vignettes:
V1 – ICU telemetry (EEG/ECG): A monitor flags “arrhythmia” during patient movement. The classifier is technically correct at times, but its rationale oscillates between lead artifact and ischemic burden–without acknowledging context. Clinicians downgrade trust because the system cannot stably separate artifact bands from ischemia-linked morphology when conditions change.
@@
V3 – Tool wear in manufacturing: Anomaly scores rise late in the wear cycle. The post-hoc explainer cites spindle torque peaks but neglects chatter harmonics that emerge earlier. Operators miss an early intervention window because explanations do not adapt to the evolving frequency structure.
@@ 2.3 Interpretability as a control objective:
  • $\Delta SP(t)=SP^{*}(t)-SP(t)$ be the interpretive error relative to a target $SP^{*}(t)$.
@@ 3 Objectives and Contributions
1. Control-theoretic formulation of interpretability
We cast interpretability as a closed-loop control problem. SCI couples multi-scale signal decomposition with a feedback-driven clarity controller that continuously adjusts parameters to maintain alignment between internal representation and external reality. Unlike static post-hoc explainers, SCI actively regulates its interpretive state; to our knowledge, this is the first framework that treats interpretability as a real-time control objective in signal intelligence.
2. Reactive multimodal decomposition
We introduce a multi-scale, multimodal representation $P(t,s)$ that captures rhythmic, trend, spatial, and cross-modal structure. Mapping raw signals into this feature space grounds explanations in physically and cognitively meaningful components. Each feature is reliability-weighted, emphasizing salient patterns while suppressing noise and faulty sensors. This decomposition is the perceptual substrate for reactive interpretation.
3. Surgical Precision (SP) as a regulated state
We define $SP(t)\in[0,1]$ as a scalar interpretive quality signal combining clarity, pattern strength, domain consistency, and predictive alignment. Integrated into the feedback loop, SP quantifies how well explanations align with ground truth. By minimizing the interpretive error $\Delta SP=SP^{*}-SP$, SCI makes interpretability a measurable, optimizable, and stabilizable control variable–analogous to how loss drives accuracy in conventional learning.
4. Closed-loop adaptation with human-in-the-loop feedback
SCI employs an online adaptation law that updates interpreter parameters $\Theta$ when $|\Delta SP|$ exceeds a threshold. Feedback sources include system discrepancies and human corrections weighted by a gain term $\lambda_{h}$, which modulates learning intensity. Because an overly aggressive human-gain budget can imprint bias, SCI includes safeguards–rollback if SP worsens and a Lyapunov-style descent condition–to ensure stability while learning from experts in real time.
5. Empirical validation across domains
We evaluate SCI on biomedical, industrial, and environmental datasets. Across these domains, SCI achieves 25–40% reductions in interpretive error (95% CI: 22%–43%), improves explanation stability (lower SP variance), and maintains predictive accuracy (AUC and F1 comparable to baselines). Ablations on feedback frequency and noise robustness indicate that the feedback-regulated equilibrium–not mere adaptation–drives these gains.
@@ 4.1 Multi-Scale Decomposition for Explainable Signal Analysis:
Classical methods decompose complex observations into analyzable primitives. Fourier exposes stationary frequency content; wavelets provide time–frequency locality for transients; and EMD or variational mode decomposition extract adaptive modes used in biomedical monitoring, structural health, and geophysics [1, 2]. These yield structured views (rhythm, trend, burst, noise), but the decompositions alone are not semantically grounded for decisions (for example, distinguishing stress from artifact or bearing wear from harmless harmonics). Recent healthcare reviews call for clinically meaningful, explainable pipelines [3], and engineering surveys urge physics-aware features in SHM [4], yet most are downstream probes–features are extracted post hoc from pretrained systems rather than formed as an interpretable substrate.
@@ 4.2 Model-Agnostic and Concept-Based XAI:
SCI addresses both. First, it ensures interpretability at the feature level via $P(t,s)$. Second, it closes the loop: explanation quality is measured as Surgical Precision $SP$, and the interpretive error $\Delta SP=SP^{*}-SP$ becomes a feedback signal that adapts interpreter parameters $\Theta$ (see §5). The interpreter $\psi_{\Theta}$ also internalizes concept-based insights through ontology-guided embeddings and concept prompts, enabling statements such as “bearing imbalance inferred from 200 Hz amplitude and sidebands” or “physiologic stress inferred from LF/HF ratio and skin-conductance reactivity,” and allowing those alignments to influence updates.
@@ 4.3 Control-Theoretic and Adaptive Interpretability:
Failure modes. We analyze robustness to erroneous or adversarial feedback and to sensor drift in §7 (ablations) and §8 (limitations).
@@ 4.4 Recent Developments (2024–2025) and Positioning:
Gap analysis and positioning. Prior art lacks a unified architecture that (i) binds semantic representation (multi-scale, multimodal, reliability-weighted), (ii) defines a scalar interpretive objective with operational meaning (our $SP$: clarity, coherence, predictive alignment), and (iii) provides a stability-aware feedback law with human-gain weighting and rollback that improves explanations without degrading predictive performance. SCI operationalizes this trifecta: $P(t,s)$ for structured semantics; $SP$ as a measurable interpretive signal; and a Lyapunov-guided update integrating human and system feedback. Empirically (§7), SCI achieves 25–40% reductions in $|\Delta SP|$ while maintaining AUC and F1 and lowering explanation variance, indicating that the system learns to sustain clarity rather than emit sporadic post hoc rationales. Conceptually, SCI reframes interpretability from a static report to a regulation problem, aligning with 2024–2025 calls for explanations that do work in the loop.
Bridge to theory. Section 6 formalizes $P(t,s)$, $\psi_{\Theta}$, $SP$ calibration, and the proof obligations for boundedness and descent of $V(\Delta SP)$ under adaptation with $\lambda_{h}$; we also specify identifiability conditions for $SP$ and analyze how reliability weights modulate gradient flow to prevent explanation drift under sensor degradation.
4.5 Guided Walk-Through: ICU ECG Lead Detachment
Concrete example (ECG lead detachment). Consider an ICU ECG where a limb lead detaches at time $t_{0}$. In (M2), $\Pi$ decomposes the raw trace into rhythmic bands (e.g., 0.5–40 Hz), low-frequency trend, and cross-sensor coherence. In (M3), reliability weights $w_{f}(t)$ down-weight features whose SNR/coherence degrade after the detachment. The interpreter $\psi_{\Theta}$ (M4) still proposes a marker (e.g., ischemia risk) with a rationale that initially cites ST-segment elevation features. The SP evaluator (M5) detects drops in $\kappa_{1}$ (clarity) and $\kappa_{3}$ (domain consistency), so $SP(t)$ falls and $\Delta SP(t)=SP^{\ast}-SP(t)$ rises. When $\lvert\Delta SP\rvert>\gamma$, the controller (M6) applies the projected update
$\Theta\leftarrow\mathrm{Proj}_{\mathcal{C}}\Big[\Theta+\eta\big(\Delta SP\nabla_{\Theta}SP+\lambda_{h}u_{h}\big)\Big].$
Within 3355 windows, the top features pivot from ischemia-like morphology to artifact-consistent bands and coherence loss; $SP(t)$ recovers toward $SP^{\star}$, and $\kappa_{3}$ returns to acceptable levels. Rollback/trust-region safeguards prevent over-correction if transient noise persists. This walk-through illustrates how SCI operationalizes interpretability as a controllable state.
@@ 5 Technical Framework
$\Delta SP(t)\triangleq SP^{*}(t)-SP(t),$
so $\Delta SP(t)>0$ when explanations lag the target and the controller should increase $SP$.
Each symbol used in this section is summarized in Table 2 for quick reference.
@@
$\Delta SP$ Interpretive error $SP^{*}-SP$
@@ 5.1 Signal Representation $P(t,s)$:
Reliability weights. SCI attaches a reliability weight $w_{f}(t)\in[0,1]$ to each feature $f\in P(t,s)$. Let
@@
Interpretability principle. $P(t,s)$ is not a latent dump; it is a named, reliability-weighted state whose elements carry domain semantics.
Notation (6.1). $X(t)$: raw inputs; $s$: spatial/structural index; $\Pi$: decomposition; $P,R,T,S,C,\Pi^{\star}$: feature blocks; $w_{f}(t)$: reliability weight.
@@ 5.2 Interpretive Mapping $\psi_{\Theta}$:
Auditability. For each $m_{k}$, SCI stores $\mathrm{TopFeat}(k,t)=\arg\max_{f}|a_{k,f}(t)|$ with signs, enabling deterministic rationales (“200 Hz line + sidebands ↑; sensor-4 temperature ↑”).
Notation (6.2). $I(t)$: interpreter output; $m_{k}$: marker; $p_{k}$: confidence; $r_{k}$: rationale; $\mathcal{D},\mathcal{V}$: priors; $\Theta$: parameters.
@@ 5.3 Interpreter, markers, and clarity $SP_{\theta}$:
where $k$ is the number of cognitive markers. In practice $g_{\theta}$ is a linear layer or shallow MLP (at most two layers), so that markers must reuse structure already present in $h_{\theta}(x)$ rather than learning an unconstrained auxiliary model.
We convert marker logits to a probability vector
@@
Here $SP_{\theta}(x)\approx 1$ when a small number of markers dominate (low entropy; a “focused” internal state) and $SP_{\theta}(x)\approx 0$ when marker usage is diffuse (high entropy). Normalizing by $\log k$ makes $SP_{\theta}$ comparable across different numbers of markers $k$.
In the streaming setting we can recover the original $SP(t)$ by aggregating per-window clarity, e.g. $SP(t)=\mathbb{E}_{x}[SP_{\theta}(x)]$ over windows ending at time $t$. We therefore treat $SP_{\theta}(x)$ as the per-example realization of the Surgical Precision signal.
5.4 Control-Theoretic and Adaptive Interpretability:
Failure modes. We analyze robustness to erroneous or adversarial feedback and to sensor drift in §7 (ablations) and §8 (limitations).
5.5 Recent Developments (2024–2025) and Positioning:
Gap analysis and positioning. Prior art lacks a unified architecture that (i) binds semantic representation (multi-scale, multimodal, reliability-weighted), (ii) defines a scalar interpretive objective with operational meaning (our $SP$: clarity, coherence, predictive alignment), and (iii) provides a stability-aware feedback law with human-gain weighting and rollback that improves explanations without degrading predictive performance. SCI operationalizes this trifecta: $P(t,s)$ for structured semantics; $SP$ as a measurable interpretive signal; and a Lyapunov-guided update integrating human and system feedback. Empirically (§7), SCI achieves 25–40% reductions in $|\Delta SP|$ while maintaining AUC and F1 and lowering explanation variance, indicating that the system learns to sustain clarity rather than emit sporadic post hoc rationales. Conceptually, SCI reframes interpretability from a static report to a regulation problem, aligning with 2024–2025 calls for explanations that do work in the loop.
Bridge to theory. Section 6 formalizes $P(t,s)$, $\psi_{\Theta}$, $SP$ calibration, and the proof obligations for boundedness and descent of $V(\Delta SP)$ under adaptation with $\lambda_{h}$; we also specify identifiability conditions for $SP$ and analyze how reliability weights modulate gradient flow to prevent explanation drift under sensor degradation.
5.6 A Guided Walk-Through of Figure 1
Concrete example (ECG lead detachment). Consider an ICU ECG where a limb lead detaches at time $t_{0}$. In (M2), $\Pi$ decomposes the raw trace into rhythmic bands (e.g., 0.5–40 Hz), low-frequency trend, and cross-sensor coherence. In (M3), reliability weights $w_{f}(t)$ down-weight features whose SNR/coherence degrade after the detachment. The interpreter $\psi_{\Theta}$ (M4) still proposes a marker (e.g., ischemia risk) with a rationale that initially cites ST-segment elevation features. The SP evaluator (M5) detects drops in $\kappa_{1}$ (clarity) and $\kappa_{3}$ (domain consistency), so $SP(t)$ falls and $\Delta SP(t)=SP^{*}-SP(t)$ rises. When $|\Delta SP|>\gamma$, the controller (M6) applies the projected update $\Theta\leftarrow\mathrm{Proj}_{\mathcal{C}}[\Theta+\eta(\Delta SP\nabla_{\Theta}SP+\lambda_{h}u_{h})]$. Within 3355 windows, the top features pivot from ischemia-like morphology to artifact-consistent bands and coherence loss; $SP(t)$ recovers. Define $\Delta SP(t)=SP^{*}(t)-SP(t)$ with time-varying target $SP^{*}(t)\in(0,1]$ (policy/ethics/physics-calibrated; specified in §3). SCI updates $\Theta$ in discrete time:

### v2

# SCI: A Metacognitive Control for Signal Dynamics

@@ Abstract
Modern deep learning systems are typically deployed as open-loop function approximators: they map inputs to outputs in a single pass, without regulating how much computation or explanatory effort is spent on a given case. In safety-critical settings this is brittle: easy and ambiguous inputs receive identical processing, and uncertainty is only read off retrospectively from raw probabilities. We introduce SCI, a closed-loop metacognitive controller layer that wraps an existing stochastic model and turns prediction into a closed-loop process. SCI monitors a scalar interpretive state $SP(t)$, here instantiated as a normalized entropy-based confidence signal, and adaptively decides whether to stop, continue sampling, or abstain. The objective is not to boost accuracy directly, but to regulate interpretive error $\Delta SP$ and expose a safety signal that tracks when the underlying model is likely to fail. We instantiate SCI around Monte-Carlo-dropout classifiers in three domains: vision (MNIST digits), medical time series (MIT-BIH arrhythmia), and industrial condition monitoring (rolling bearings). In all cases, the controller allocates more inference steps to misclassified inputs than to correct ones (up to $\approx 3$$4\times$ on MNIST and bearings, and $1.4\times$ on MIT-BIH). The resulting $\Delta SP$ acts as a usable safety signal for detecting misclassifications (AUROC $0.63$ on MNIST, $0.70$ on MIT-BIH, $0.86$ on bearings).
Code and reproducibility: https://github.com/vishal-1344/sci
@@ 1 Introduction
In safety-critical settings, the usefulness of an alert depends as much on its rationale as on its score. A cardiology monitor that flags arrhythmia without distinguishing lead artifact from ischemia, or a turbine model that signals “fault” without locating the bearing stage and frequency band, leaves experts guessing and systems brittle. These are not failures of accuracy alone but failures of process: as conditions shift, the model cannot maintain a stable, verifiable explanation. By actively regulating its own interpretive state, SCI endows standard models with a form of metacognitive control: the ability to recognize ambiguity and allocate computational resources to resolve it autonomously.
@@
SCI was conceived to close this gap. Earlier iterations laid the groundwork. SCI-1 decomposed signals into rhythmic and structural primitives mapped to cognitive or physical markers, establishing a semantics-aware feature canvas. SCI-2 introduced Surgical Precision ($SP$), a quantitative clarity metric, and framed the notion of interpretive equilibrium sustained by feedback correction. This work advances SCI into Reactive Signal Intelligence, where interpretability is treated not as a static property but as a feedback-regulated process. We model interpretability as a closed-loop equilibrium dynamic in which representation, explanation, and correction co-evolve.
Within this control-theoretic formulation, interpretation becomes a quantifiable control objective: the system continuously minimizes an interpretive discrepancy $\Delta SP$ to align internal reasoning with domain reality. SCI instantiates this dynamic view through three integrated components:
@@
  • A Lyapunov-guided adaptive controller with a human-feedback gain budget $\lambda_{h}$, descent conditions, and stability safeguards (with $\lambda_{h}<\mu/(Uc)$ and trust-region/rollback, $V=\tfrac{1}{2}(\Delta SP)^{2}$ decreases monotonically up to bounded noise; see §5.4).
This reframes transparency as a control problem that sustains equilibrium, maintaining both performance and clarity under sensor drift, nonstationarity, and bounded perturbations (adversarial or human).
@@
  • Stability and Evidence. We provide theoretical stability conditions and empirical validation across vision, medical, and industrial domains. Results demonstrate emergent metacognition (allocating $3.6\times$$3.8\times$ more compute to ambiguous inputs), validate $\Delta SP$ as a safety signal with AUROC $\approx 0.70$$0.86$, and confirm that the controller matches the accuracy of fixed ensembles with greater efficiency.
@@
Section 2 motivates reactivity with domain examples. Section 3 states objectives and delineates contributions. Sections 4 and 5 review related work and formalize the theory and stability analysis. Section 6 details the architecture and feedback mechanisms. Section 7 instantiates SCI on three public benchmarks (MNIST digits, MIT-BIH ECG, and IMS/NASA bearing vibration) and characterizes its behavior as a metacognitive controller. Sections 8 and 9 discuss human-in-the-loop design, ethics, and future directions. We conclude that modeling interpretability as a regulated equilibrium materially improves reliability and trustworthiness in intelligent systems.
@@
SCI in 60 seconds. What: Treat explanation quality as a signal $SP(t)$ with target $SP^{\star}$. How: When $|\Delta SP|{=}|SP^{\star}-SP|$ is large, update $\Theta$ via $\Theta\leftarrow\mathrm{Proj}_{\mathcal{C}}\!\big[\Theta+\eta(\Delta SP\,\nabla_{\Theta}SP+\lambda_{h}u_{h})\big]$, bounded by rollback/trust-region and a human-gain budget $\lambda_{h}<\mu/(Uc)$. Why it works: With these bounds, the Lyapunov energy $V=\tfrac{1}{2}(\Delta SP)^{2}$ decreases (up to noise), stabilizing rationales while preserving task accuracy. Payoff: In Vision, Medical, and Industrial tasks, SCI autonomously allocates $3.6\times$$3.8\times$ more compute to errors than correct predictions and provides a safety signal ($\Delta SP$) with AUROC 0.700.700.860.86.
@@ 2.1 Real-world failure vignettes:
V1 – ICU telemetry (ECG/EEG, illustrative): For example, a bedside monitor may flag “arrhythmia” during patient movement. The classifier can be technically correct at times, but its rationale oscillates between lead artifact and ischemic burden—without acknowledging context. Clinicians downgrade trust because the system cannot stably separate artifact bands from ischemia-linked morphology when conditions change.
@@
V3 – Tool wear in manufacturing (illustrative): In a late wear cycle, anomaly scores may rise only near the end. A post-hoc explainer might cite spindle torque peaks but neglect chatter harmonics that emerge earlier. Operators can miss an early intervention window because explanations do not adapt to the evolving frequency structure.
@@ 2.3 Interpretability as a control objective:
  • $\Delta SP(t)=SP^{\star}(t)-SP(t)$ be the interpretive error relative to a target $SP^{\star}(t)$.
@@ 3 Objectives and Contributions
1. Control-theoretic formulation of interpretability. We cast interpretability as a closed-loop control problem. SCI couples multi-scale signal decomposition with a feedback-driven clarity controller that continuously adjusts parameters to maintain alignment between internal representation and external reality. Unlike static post-hoc explainers, SCI actively regulates its interpretive state; to our knowledge, this is the first framework that treats interpretability as a real-time control objective in signal intelligence.
2. Reactive multimodal decomposition. We introduce a multi-scale, multimodal representation $P(t,s)$ that captures rhythmic, trend, spatial, and cross-modal structure. Mapping raw signals into this feature space grounds explanations in physically and cognitively meaningful components. Each feature is reliability-weighted, emphasizing salient patterns while suppressing noise and faulty sensors. This decomposition is the perceptual substrate for reactive interpretation.
3. Surgical Precision (SP) as a regulated state. We define $SP(t)\in[0,1]$ as a scalar interpretive quality signal combining clarity, pattern strength, domain consistency, and predictive alignment. Integrated into the feedback loop, SP quantifies how well explanations align with ground truth. By minimizing the interpretive error $\Delta SP=SP^{\star}-SP$, SCI makes interpretability a measurable, optimizable, and stabilizable control variable, analogous to how loss drives accuracy in conventional learning.
4. Closed-loop adaptation with human-in-the-loop feedback. SCI employs an online adaptation law that updates interpreter parameters $\Theta$ when $|\Delta SP|$ exceeds a threshold. Feedback sources include system discrepancies and human corrections weighted by a gain term $\lambda_{h}$, which modulates learning intensity. Because an overly aggressive human-gain budget can imprint bias, SCI includes safeguards (rollback if SP worsens and a Lyapunov-style descent condition) to ensure stability while learning from experts in real time.
5. Empirical validation across domains. We evaluate SCI on biomedical (MIT-BIH), industrial (Bearings), and vision (MNIST) benchmarks. Across these settings, the controller (i) allocates significantly more computation to hard inputs ($3.6\times$$3.8\times$ cognitive ratio), (ii) produces an interpretive error signal $\Delta SP$ that detects misclassifications with AUROC $\approx 0.70$$0.86$, and (iii) matches the accuracy of fixed-depth ensembles on ECG while using fewer samples on average.
@@ 4.1 Multi-Scale Decomposition for Explainable Signal Analysis:
Classical methods decompose complex observations into analyzable primitives. Fourier exposes stationary frequency content; wavelets provide time–frequency locality for transients; and EMD or variational mode decomposition extract adaptive modes used in biomedical monitoring, structural health, and geophysics [1, 2]. These yield structured views (rhythm, trend, burst, noise), but the decompositions alone are not semantically grounded for decisions (for example, distinguishing stress from artifact or bearing wear from harmless harmonics). Recent healthcare reviews call for clinically meaningful, explainable pipelines [3], and engineering surveys urge physics-aware features in SHM [4], yet most are downstream probes where features are extracted post hoc from pretrained systems rather than formed as an interpretable substrate.
@@ 4.2 Model-Agnostic and Concept-Based XAI:
SCI addresses both. First, it ensures interpretability at the feature level via $P(t,s)$. Second, it closes the loop: explanation quality is measured as Surgical Precision $SP$, and the interpretive error $\Delta SP=SP^{\star}-SP$ becomes a feedback signal that adapts interpreter parameters $\Theta$ (see §5). The interpreter $\psi_{\Theta}$ also internalizes concept-based insights through ontology-guided embeddings and concept prompts, enabling statements such as “bearing imbalance inferred from 200 Hz amplitude and sidebands” or “physiologic stress inferred from LF/HF ratio and skin-conductance reactivity,” and allowing those alignments to influence updates.
@@ 4.4 Recent Developments (2024–2025) and Positioning:
Gap analysis and positioning.
Prior art lacks a unified architecture that (i) binds semantic representation, (ii) defines a scalar interpretive objective, and (iii) provides a stability-aware feedback law. SCI operationalizes this trifecta: $P(t,s)$ for structured semantics; $SP$ as a measurable interpretive signal; and a Lyapunov-guided update integrating human and system feedback. Empirically (§7), SCI demonstrates emergent metacognition (allocating $3.6\times$ more compute to hard inputs) and validates $\Delta SP$ as a robust safety signal (AUROC 0.700.700.860.86), reframing interpretability from a static report to a regulated equilibrium process.
Bridge to theory.
Section 6 formalizes $P(t,s)$, $\psi_{\Theta}$, $SP$ calibration, and the proof obligations for boundedness and descent of $V(\Delta SP)$ under adaptation with $\lambda_{h}$; we also specify identifiability conditions for $SP$ and analyze how reliability weights modulate gradient flow to prevent explanation drift under sensor degradation.
@@ 5 Technical Framework
$\Delta SP(t)\triangleq SP^{\star}(t)-SP(t),$
so $\Delta SP(t)>0$ when explanations lag the target and the controller should increase $SP$. Each symbol used in this section is summarized in Table 2 for quick reference.
@@
$\Delta SP$ Interpretive error $SP^{\star}-SP$
@@ 5.1 Signal Representation $P(t,s)$:
Reliability weights.
SCI attaches a reliability weight $w_{f}(t)\in[0,1]$ to each feature $f\in P(t,s)$. Let
@@
Interpretability principle.
$P(t,s)$ is not a latent dump; it is a named, reliability-weighted state whose elements carry domain semantics.
@@ 5.2 Interpretive Mapping $\psi_{\Theta}$:
Auditability.
For each $m_{k}$, SCI stores $\mathrm{TopFeat}(k,t)=\arg\max_{f}|a_{k,f}(t)|$ with signs, enabling deterministic rationales (“200 Hz line + sidebands ↑; sensor-4 temperature ↑”).
@@ 5.3 Interpreter, markers, and clarity $SP_{\theta}$:
where $k$ is the number of cognitive markers. In practice $g_{\theta}$ is a linear layer or shallow MLP (at most two layers), so that markers must reuse structure already present in $h_{\theta}(x)$ rather than learning an unconstrained auxiliary model. We convert marker logits to a probability vector
@@
Here $SP_{\theta}(x)\approx 1$ when a small number of markers dominate (low entropy; a “focused” internal state) and $SP_{\theta}(x)\approx 0$ when marker usage is diffuse (high entropy). Normalizing by $\log k$ makes $SP_{\theta}$ comparable across different numbers of markers $k$. In the streaming setting we can recover the original $SP(t)$ by aggregating per-window clarity, e.g. $SP(t)=\mathbb{E}_{x}[SP_{\theta}(x)]$ over windows ending at time $t$. We therefore treat $SP_{\theta}(x)$ as the per-example realization of the Surgical Precision signal.
5.4 Closed-Loop Update and Lyapunov Energy
Define $\Delta SP(t)=SP^{\star}(t)-SP(t)$ with time-varying target $SP^{\star}(t)\in(0,1]$ (policy/ethics/physics-calibrated; specified in §3). SCI updates $\Theta$ in discrete time: